### v1

# Symmetric Entropy-Constrained Video Coding for Machines

Abstract
As video transmission increasingly serves machine vision systems (MVS) instead of human vision systems (HVS), video coding for machines (VCM) has become a critical research topic. Existing VCM methods often bind codecs to specific downstream models, requiring retraining or supervised data and thus limiting generalization in multi-task scenarios. Recently, unified VCM frameworks have employed visual backbones (VB) and visual foundation models (VFM) to support multiple video understanding tasks with a single codec. They mainly utilize VB/VFM to maintain semantic consistency or suppress non-semantic information, but seldom explore how to directly link video coding with understanding under VB/VFM guidance. Hence, we propose a Symmetric Entropy-Constrained Video Coding framework for Machines (SEC-VCM). It establishes a symmetric alignment between the video codec and VB, allowing the codec to leverage VB’s representation capabilities to preserve semantics and discard MVS-irrelevant information. Specifically, a bi-directional entropy-constraint (BiEC) mechanism ensures symmetry between the process of video decoding and VB encoding by suppressing conditional entropy. This helps the codec to explicitly handle semantic information beneficial for MVS while squeezing useless information. Furthermore, a semantic–pixel dual-path fusion (SPDF) module injects pixel-level priors into the final reconstruction. Through semantic-pixel fusion, it suppresses artifacts harmful to MVS and improves machine-oriented reconstruction quality. Experimental results show our framework achieves state-of-the-art (SOTA) in rate–task performance, with significant bitrate savings over VTM on video instance segmentation (37.41%), video object segmentation (29.83%), object detection (46.22%), and multiple object tracking (44.94%). We will release our code.
I Introduction
Figure 1: Overview of different methods. $I(\cdot)$ denotes mutual information, $H(\cdot)$ denotes Shannon entropy, and $\Uparrow$” denotes that the optimization goal is to maximize the value. $\mathcal{E}_{\textit{visual}}$ denotes the encoder of visual backbone (VB) or visual foundation model (VFM). (a) Standard or neural video codecs [1, 2] are optimized for HVS, which makes the decoded videos $\hat{x}$ suboptimal for machine vision. (b) Each neural video/feature codec [3, 4] is optimized for a specific machine vision task. (c) The unified codec [5, 6, 7, 8] is optimized with the assistance of the VB or VFM to reconstruct videos suitable for multiple downstream tasks. (d) Our unified compression framework aligns the video coding process with the video understanding process from the entropy perspective and squeezes machine-vision-irrelevant information during the process of video reconstruction.
Digital video plays a crucial role in our daily lives, comprising a significant portion of data storage. To efficiently store and transmit this massive amount of data, video coding techniques have been developed, primarily designed to meet the requirements of the human visual system (HVS). For videos aimed at HVS, such as movies and short clips, it is essential to preserve visual details perceptible to humans during compression. Meanwhile, videos collected for machine vision tasks, such as surveillance analysis [9] and facial recognition [10], typically require only part of the video information, rather than the complete visual details [11, 12]. Currently, neural video coding frameworks have evolved significantly and offer excellent compression performance [13, 2, 14, 15, 16]. However, these methods are optimized primarily for HVS-related metrics and do not fully satisfy the requirements of machine vision systems (MVS) in terms of video understanding (e.g., detection, segmentation, and tracking) [17], as shown in Fig. 1(a). Therefore, it is essential to develop video coding for machines (VCM) methods and bridge the gap between video coding and video understanding.
To reconstruct machine-vision-friendly videos, most existing methods [18, 19, 4] bridge the upstream codec with specific downstream task models and jointly optimize with the loss functions of different tasks, as shown in Fig. 1(b). However, this requires retraining the entire upstream codec or downstream models for each new task, resulting in significant computational and time costs. Meanwhile, some parameter-efficient fine-tuning (PEFT) strategies are introduced to adapt codecs to new downstream tasks with lower cost [3]. Nonetheless, most of them still rely on the labeled data and supervised training process for specific downstream tasks, limiting their generalization in multi-task scenario.
To improve the applicability, some methods [20, 8] have adopted unified compression frameworks to reconstruct videos suitable for multiple downstream tasks simultaneously. These methods abandon supervised training and instead maintain perceptual consistency in a self-supervised manner. Since the visual backbone (VB) and visual foundation model (VFM), such as ResNet50 [21], DINOv2 [22], and Mask Auto-Encoder (MAE) [23], are capable of extracting general feature representations for video understanding [20], they are introduced to maintain semantic consistency between the original and reconstructed frames [5, 6, 7, 24], as shown in Fig. 1(c). Additionally, some methods [5, 6] have extended an analysis-oriented bitstream beyond standard codecs (e.g., H.264/AVC [25] and H.266/VVC [26]) and introduced VB and VFM to suppress non-semantic information in the analysis-oriented bitstream. However, these approaches still treat VB/VFM merely as auxiliary tools for enforcing semantic consistency or non-semantic suppressing, without fully exploiting their potential to guide the video codec itself. This limitation leaves the video coding process loosely connected to the video understanding process by machines, motivating us to design a framework that explicitly aligns the two.
To address the above issues, we propose a Symmetric Entropy-Constrained Video Coding framework for Machine (SEC-VCM). It supports multiple downstream tasks with a single unsupervised training process and further eliminates machine-vision-irrelevant information by leveraging existing visual backbones, as shown in Fig. 1(d). Its key principle is to bridge the gap between the video coding and understanding processes from an entropy perspective, explicitly guiding the video codec to handle only information beneficial for video understanding while discarding irrelevant information. Specifically, an information-alignment-based semantic decoder with a bi-directional entropy-constraint (BiEC) mechanism is proposed to make the video decoding process of the codec symmetric and consistent with the encoding process of the visual backbone. In addition, a semantic-pixel dual-path fusion (SPDF) module is designed to incorporate HVS-related priors into the machine-oriented reconstruction process. It helps reduce artifacts harmful to MVS and fulfill the frame content through a gated fusion mechanism. The experiments on four machine vision tasks demonstrate that the proposed SEC-VCM achieves state-of-the-art (SOTA) rate-task performance. Compared to VTM, our method achieves significant bitrate reductions, with savings of 37.41% for video instance segmentation, 29.83% for video object segmentation, 46.22% for object detection, and 44.94% for multiple object tracking.
Our main contributions are summarized as follows:
  • We propose SEC-VCM, a novel video coding framework explicitly aligned with machine vision requirements from an information-theoretic perspective, enabling general support for diverse downstream machine tasks.
  • We introduce an information-alignment-based semantic decoder with a bi-directional entropy-constraint mechanism, which aligns the codec’s decoding process with the visual backbone’s encoding process, thereby eliminating machine-vision-irrelevant redundancy.
  • We design a semantic-pixel dual-path fusion module to inject pixel-level prior knowledge into the machine-oriented decoding process for artifact suppression, details fulfillment, and improved downstream tasks accuracy.
II Related Works
II-A Neural Video Coding
Most existing neural video coding methods [27, 28, 29, 14, 30, 31, 32, 15, 16, 33, 34] follow the motion-then-residual paradigm and mainly include inter-prediction and residual/context compression. Lu et al. [27] proposed the first end-to-end video coding framework DVC, which uses the optical flow [35] for inter-prediction. Agustsson et al. [30] proposed the Scale Space Flow and added a Gaussian dimension to create more reference pixels for inter-frame prediction. Hu et al. [36] proposed FVC to convert the video from the pixel domain to the feature domain and used deformable convolution for ME-MC in the feature domain. Li et al. [28] improved the concept of predictive coding and proposed the conditional coding paradigm for better compression performance, termed DCVC. Sheng et al. [37] upgraded the DCVC and proposed DCVC-TCM with a multi-scale conditional coding to better remove temporal redundancy. Furthermore, many methods [15, 2, 38] introduce long-range training to suppress accumulation error and make the codec perform better on a long group of pictures (GOP). Overall, existing neural video coding methods improve compression efficiency from various perspectives, such as inter- or bi-directional prediction [30, 15, 14, 16, 39, 40], residual/context compression [13], and entropy models [29, 13]. Many NVC methods, such as ECVC-series [38] and DCVC-series [14, 2, 13, 15, 39], demonstrate excellent compression performance.
II-B Neural Video Coding for Machines
The field of video coding for machines mainly includes two paradigms: analyze-then-compress (ATC) and compress-then-analyze (CTA). The ATC paradigm focuses on feature coding, while the CTA paradigm focuses on machine-vision-oriented image/video coding.
II-B1 Analyze-then-Compress Paradigm
Since many edge devices (e.g., Smartphones, AI cameras) support neural processing, they can extract features from images or videos, compress them, and transmit these features to reduce the computational load on the server. Most methods rely on labeled data and joint optimization strategies to make neural feature codecs meet the needs of specific downstream tasks [41, 42, 18, 19, 43]. However, these methods rely on retraining the upstream codec or downstream task models to adapt to newly coming downstream tasks. Some works introduced lightweight adapters for latent-domain [44] or feature-domain [17] transformation, enabling the reconstructed features to satisfy multiple downstream tasks and avoiding heavy fine-tuning.
Figure 2: Overview of the proposed SEC-VCM framework. (1) The basic compression loop contains motion and contextual coding processes. “Q” denotes the quantization operation. Notably, some modules in this step are not shown above for simplicity, since this step is directly adopted from the neural video codec DCVC-HEM [29]. (2) The semantic decoding process is information-aligned with the encoding process of the pretrained visual backbone by the proposed BiEC mechanism from an entropy perspective. This helps the semantic decoder $\mathcal{D}_{\textit{sem}}$ purely handle information beneficial for MVS. (3) The pixel-level prior knowledge $f_{\textit{pixel}}$ is injected into the perception-oriented decoding process to fulfill frame content, suppress artifacts, and enhance final reconstruction quality.
II-B2 Compress-then-Analyze Paradigm
Some CTA methods cascade neural codecs or pre-process modules with downstream tasks and adopt jointly supervised training to achieve promising rate-task performance on specific tasks (e.g., object detection, face recognition, optical character recognition) [45, 46, 47, 48]. To reduce training costs, some methods explored parameter-efficient fine-tuning (PEFT) based strategies of the upstream codecs for specific downstream tasks [3, 49]. However, these methods need to fine-tune the entire codec or specific modules for each downstream task, which increases costs and limits their generalization. Meanwhile, some methods abandoned task-specific joint training and proposed unified VCM codecs generalizable to multiple downstream tasks. Tian et al. proposed to enhance edge [20] and semantic information [5, 6] by introducing an additional bitstream beyond the standard codecs. Other methods incorporated the visual backbone (VB) or visual foundation model (VFM) for self-supervised training. Shindo et al. proposed the SA-ICM [7], which uses the Segment Anything Model (SAM) to constrain the semantic and edge consistency between the original and reconstructed images. Tian et al. [8] employed DINO-v2 [22] and Video-MAE [23] in the latent domain to maintain semantic integrity. These methods typically apply VB and VFM to compute feature-domain losses for semantic consistency, yet they have not fully explored the alignment of video coding and video understanding at the information level, nor investigated how video understanding can explicitly guide the video coding process.
III Methodology
III-A Motivation and Overview
Existing neural video codecs typically adopt VAE-like networks [50] to compress the input video $x$ into compact latent representations and reconstruct the frame $\hat{x}$ for the machine vision system (MVS). Current methods optimize the codec for specific or multiple machine vision tasks, either in a supervised or unsupervised manner [4, 43, 5, 6]. Also, some methods maintain semantic consistency by minimizing the perceptual distance between the original video $x$ and compressed video $\hat{x}$ measured by visual backbone (VB) or visual foundation model (VFM) [8, 7]. Their paradigm is formulated as follows:
[…]
where $\mathcal{E}(\cdot)$ and $\mathcal{D}(\cdot)$ denote the encoder and decoder of the video codec, $\lfloor\cdot\rceil$” denotes the quantization operation. $\mathcal{L}$ denotes the total loss, and $R$ denotes the bitrate. $\mathcal{E}_{\textit{visual}}$ denotes the encoder of VB or VFM. $D(\cdot,\cdot)$ denotes pixel-domain or feature-domain distortion (e.g., MSE).
Although VB and VFM have been introduced to help codecs preserve semantics, their integration with video coding remains insufficiently explored. They excel at extracting rich and general feature representations for multiple machine vision tasks [21, 22], yet such features are not naturally suited for entropy coding or direct compression. To better leverage their general capacity, we align the video codec with VB/VFM from an entropy perspective, encouraging it to selectively preserve features beneficial for MVS while squeezing irrelevant information. The alignment objective is formulated as:
$\mathcal{D}_{\textit{sem}}(\hat{y})\xleftrightarrow{\text{aligned}}\mathcal{E}_{\textit{visual}}(x)\xrightarrow{\text{capable}}\text{downstream tasks},$ (2)
where $\mathcal{D}_{\textit{sem}}(\cdot)$ denotes the semantic decoder and $\hat{y}$ is the transmitted bitstream (entropy-coded latent representation). Additionally, completing the compression loop requires low-level details to support inter-prediction [4]. Therefore, we propose that pixel-oriented reconstruction and machine-oriented reconstruction are conducted separately, with the former providing pixel-level priors for the latter.
Figure 3: Structures of proposed modules. (a) Structure of the information-alignment-based semantic decoder and BiEC mechanism. DEDE denotes the distribution estimation module, and $\uparrow\times 2$” denotes two times up-sampling. During training, the intermediate features for the decoding $f_{\textit{MVS}}$ are made hierarchically symmetric with the features $f_{\textit{VB}}$ from the visual backbone. (b) Structure of the SPDF module. Machine vision system (MVS)-oriented semantic feature $f_{t}^{\textit{MVS}}$ and pixel-related feature $f_{t}^{\textit{pixel}}$ are selectively combined based on the factor $\alpha$.
To achieve the above ideas, we propose the Symmetric Entropy-Constrained Video Coding framework for Machine (SEC-VCM), as shown in Fig. 2. Specifically, it consists of three main steps: inter-prediction and contextual coding for the basic compression loop, machine-oriented semantic decoding with the Bi-directional Entropy-Constraint (BiEC) mechanism, and semantic-pixel gated fusion. (1) In the first step, we inherit the basic compression loop from DCVC-HEM [29], where motion and contextual information are compressed, quantized, entropy-coded, and then reconstructed. The compensated features $c_{1}$, $c_{2}$, and $c_{3}$ serve as motion priors for both contextual coding and our semantic decoding. The motion latent $\hat{y}_{t}^{mv}$ and contextual latent $\hat{y}_{t}^{ctx}$ are then transmitted in the bitstream. (2) In the second step, the contextual latent representation $\hat{y}t^{ctx}$ is employed for semantic decoding, where the intermediate feature $f_{\textit{MVS}}$ for MVS-oriented reconstruction is symmetrically aligned with $f_{\textit{VB}}$ extracted by the visual backbone $\mathcal{E}_{\textit{visual}}$. Specifically, in the BiEC mechanism, the symmetry is enforced at three scales, and the alignment is achieved by constraining conditional entropy. (3) In the third step, since $f_{\textit{MVS}}$ mainly contains the semantic information, additional detailed pixel-level priors $f_{\textit{pixel}}$ are injected into the final reconstruction process to enrich frame details, suppress artifacts, and produce frames better suited for MVS. Notably, our framework is general across multiple downstream tasks, and we therefore remove the task loss $\mathcal{L}_{task(\hat{x})}$ that relies on specific labeled datasets. Our paradigm is illustrated as follows:
$\left\{\begin{aligned} \hat{x}=&\mathcal{D}_{\textit{sem}}(\lfloor\mathcal{E}(x)\rceil|\mathcal{E}_{\textit{visual}}(x),f_{\textit{pixel}}),\\ \mathcal{L}=&R+D(x,\hat{x})+D(\mathcal{E}_{\textit{visual}}(x),\mathcal{E}_{\textit{visual}}(\hat{x}))\\ &+H(f_{\textit{MVS}}|f_{\textit{VB}})+H(f_{\textit{VB}}|f_{\textit{MVS}}),\end{aligned}\right$. (3)
where $\mathcal{D}_{\textit{sem}}(\cdot)$ denotes the proposed information-alignment-based semantic decoder, $f_{\textit{VB}}$ denotes the feature extracted by the visual backbone, $f_{\textit{MVS}}$ denotes the feature in the semantic decoder, and $f_{\textit{pixel}}$ denotes the feature for low-level reconstruction. $H(f_{\textit{MVS}}|f_{\textit{VB}})$ and $H(f_{\textit{VB}}|f_{\textit{MVS}})$ denote the conditional entropy terms used as loss items in the proposed BiEC mechanism.
III-B Semantic Decoding with Symmetric Entropy-Constraint
To leverage the visual backbone’s strong capability of extracting machine-vision-oriented feature representations, the video decoding process is hierarchically and symmetrically aligned with the encoding process of VB during training. It enables the decoder to inherit the prior knowledge learned by the backbone that has experienced large-scale pretraining, focusing purely on features relevant to machine vision tasks. To achieve this idea, an information-aligned semantic decoder and a BiEC mechanism are designed, as shown in Fig. 3(a).
III-B1 Information-Alignment-based Semantic Decoder
The semantic decoder comprises four bottleneck residual blocks and three up-sampling layers. It progressively recovers the quantized and compact latent representation $\hat{y}_{ctx}$ (at $1/16$ resolution) back to the original spatial resolution, producing machine-vision-oriented features $f_{\textit{MVS}}$. The inter-predicted features $c_{2}$ ($1/2$ size) and $c_{3}$ ($1/4$ size) are used as conditions for frame reconstruction, referring to [28, 29]. Notably, intermediate features $f_{\textit{MVS}}$ within the semantic decoder are aligned with the intermediate results $f_{\textit{VB}}$ from the visual backbone at three distinct scales. These three hierarchical alignments make the decoder symmetric to the pretrained guidance backbone.
III-B2 Bi-directional Entropy-Constraint Mechanism
The features $f_{\textit{VB}}$ extracted by the pretrained backbone are optimized for analysis but not directly suitable for frame reconstruction. Therefore, classic alignment metrics such as MSE or KL-divergence are ineffective when applied between $f_{\textit{VB}}$ and the feature $f_{\textit{MVS}}$ in the video decoding process. Instead, we propose to align $f_{\textit{VB}}$ and the reconstruction-oriented feature $f_{\textit{MVS}}$ from an information-theoretic perspective. The core idea is to ensure that both feature sets carry equivalent information, allowing accurate inference of one from the other. The distribution estimation (DE) module is introduced to estimate the element-wise distribution parameters $(\mu_{\textit{MVS}},\sigma_{\textit{MVS}})$ of $f_{\textit{MVS}}$ based on $f_{\textit{VB}}$. A more accurate parameter estimation implies a higher conditional probability and thus a lower conditional entropy, indicating that $f_{\textit{VB}}$ retains all information necessary to infer $f_{\textit{MVS}}$, therefore $f_{\textit{MVS}}$ is a subset of $f_{\textit{VB}}$. The above process is formulated as follows:
$\mu_{\textit{MVS}}^{i},\sigma_{\textit{MVS}}^{i}$ $=DE(f_{\textit{VB}}^{i}),$ (4)
$H(f_{\textit{MVS}}|f_{\textit{VB}})\Downarrow$ $\Rightarrow\sum_{i}p(f^{i}_{\textit{MVS}}|\mu^{i}_{\textit{MVS}},\sigma^{i}_{\textit{MVS}})\Uparrow,$
where $\mu_{\textit{MVS}}^{i}$ and $\sigma_{\textit{MVS}}^{i}$ denote estimated element-wise distribution parameters of $f_{\textit{MVS}}^{i}$. $i\in\{1/4,1/8,1/16\}$ denotes the layer-index of $f_{\textit{MVS}}$, $DE(\cdot)$ denotes the distribution estimation module, $H(\cdot|\cdot)$ denotes the conditional entropy, and $p(\cdot|\cdot)$ denotes conditional probability [50]. $\Rightarrow$” denotes the transformation of the problem definition. $\Downarrow$” and $\Uparrow$” denote the optimization goal, minimizing the conditional entropy and maximizing the conditional probability, respectively.
Figure 4: The process of the BiEC mechanism. Information alignment between $f_{\textit{MVS}}$ for decoding and $f_{\textit{VB}}$ from the visual backbone is achieved through the constraint of conditional entropy in two distinct directions. $\Downarrow$” indicates that the optimization goal is to minimize the item.
Figure 5: The t-SNE visualization of intermediate features $f_{\textit{MVS}}^{1/8}$ in semantic decoder with and without symmetric information alignment. After alignment via the BiEC mechanism, intra-class distances are effectively reduced, and the intermediate features $f_{\textit{MVS}}^{1/8}$ in the decoding process exhibit more discriminative semantic representations.
To enforce symmetry, we further impose a conditional entropy constraint in the reverse direction, making $f_{\textit{VB}}$ a subset of $f_{\textit{MVS}}$. Through this bi-directional constraint, the features $f_{\textit{MVS}}$ in the semantic decoder and the visual-backbone-extracted representations $f_{\textit{VB}}$ are softly aligned, as illustrated in Fig. 4. The above process is formulated as follows:
$\mu_{\textit{VB}}^{i},\sigma_{\textit{VB}}^{i}$ $=DE(f_{\textit{MVS}}^{i}),$ (5)
$H(f_{\textit{VB}}|f_{\textit{MVS}})\Downarrow$ $\Rightarrow\sum_{i}p(f^{i}_{\textit{VB}}|\mu^{i}_{\textit{VB}},\sigma^{i}_{\textit{VB}})\Uparrow,$
where $\mu_{\textit{VB}}^{i}$ and $\sigma_{\textit{VB}}^{i}$ denote estimated element-wise distribution parameters of $f_{\textit{VB}}^{i}$.
Specifically, each distribution estimation module contains a convolution layer and a ResBlock [21]. Estimated distribution parameters $\mu$ and $\sigma$ have the same shape as the corresponding feature $f$. Following existing methods [50, 29], the Laplace distribution is used to model the conditional probability of the features $f_{\textit{MVS}}$ and $f_{\textit{VB}}$. The conditional entropy $H(f_{\textit{VB}}|f_{\textit{MVS}})$ and $H(f_{\textit{MVS}}|f_{\textit{VB}})$ are calculated as follows:
$H(f_{\textit{VB}}|f_{\textit{MVS}})$ $=\sum_{i}\sum_{j}-log_{2}(p(f_{\textit{VB}}^{i,j}|\mu_{\textit{VB}}^{i,j},\sigma_{\textit{VB}}^{i,j})),$ (6)
$H(f_{\textit{MVS}}|f_{\textit{VB}})$ $=\sum_{i}\sum_{j}-log_{2}(p(f_{\textit{MVS}}^{i,j}|\mu_{\textit{MVS}}^{i,j},\sigma_{\textit{MVS}}^{i,j})),$
where $i$ and $j$ denote the layer and position indices, respectively.
Figure 6: Visualization results of reconstructed frame using different feature types at $bpp\approx 0.08$. Relying only on pixel-oriented feature $f_{\textit{pixel}}$ tends to miss fine details of the dog’s fur textures, while using only semantic-level feature $f_{\textit{semantic}}$ leads to the loss of grass texture and introduces artifacts. By fusing $f_{\textit{semantic}}$ and $f_{\textit{pixel}}$, our method reconstructs the dog’s fur details better and avoids artifacts in the grass.
III-B3 Discussion
Through the BiEC mechanism, the semantic decoder is encouraged to reconstruct semantic representations by relying only on features that are beneficial for video understanding, guided by the visual backbone. In other words, the intermediate features $f_{\textit{MVS}}$ of the semantic decoder are expected to be semantic-aware. To illustrate this property, we provide a visualization. Specifically, 400 images from 8 classes of the ImageNet-1k [21] validation set are fed into our framework, and the intermediate features $f_{\textit{MVS}}^{1/8}$ of the semantic decoder are projected into a two-dimensional space using t-SNE [51]. Since our framework is designed for P-frames and cannot be evaluated on a single frame, the reference frame is set to all zeros for this visualization. As shown in Fig. 5, the semantic decoder, under the joint effect of the visual backbone and the BiEC mechanism, exhibits improved capability in forming compact and semantically meaningful feature representations.
III-C Semantic-Pixel Dual-Path Fusion
The BiEC-based semantic decoder focuses on features relevant to machine vision tasks, while the motion- and context-related modules aim to reconstruct pixel-level details. These two objectives are not fully aligned. Although machine-oriented tasks and traditional distortion metrics (e.g., PSNR, MS-SSIM) both emphasize structural and color preservation, they differ in their sensitivity to low-level textures and patterns. To bridge this gap, we propose the SPDF module, which injects pixel-level priors into the machine-oriented reconstruction by selectively combining pixel-level features $f_{\textit{pixel}}$ and semantic-level features $f_{\textit{MVS}}$. Specifically, as illustrated in Fig. 6, using only $f_{\textit{MVS}}$ helps retain the texture of the dog’s fur and may increase its detection accuracy, but it ignores the texture of the grass and introduces visible artifacts. In contrast, relying solely on $f_{\textit{pixel}}$ leads to blur under limited bitrates. By gated fusing both $f_{\textit{MVS}}$ and $f_{\textit{pixel}}$, our method effectively suppresses grass artifacts and preserves fine textures of fur.
The structure of the SPDF module is shown in Fig. 3(b). Specifically, the machine-oriented feature $f_{\textit{MVS}}$ and pixel-level features $f_{\textit{pixel}}$ are both fed into the SPDF module. A residual block is employed to calculate an element-wise gating coefficient $\alpha$, which is formulated as follows:
$\alpha$ $=Sigmoid(ResBlock(Concat(f_{\textit{MVS}},f_{\textit{pixel}}))),$ (7)
where $ResBlock(\cdot)$ denotes Bottleneck-ResBlock [21], $Concat(\cdot,\cdot)$ denotes channel-wise concatenation. Two stacked U-Nets are employed to perform deeper feature transformations on $f_{\textit{MVS}}$ and get the semantic-level feature $f_{\textit{semantic}}$. Subsequently, $f_{\textit{semantic}}$ and $f_{\textit{pixel}}$ are linearly combined based on the coefficient $\alpha$. After fusion, another U-Net is used to refine the combined features, and a Conv2D layer is applied to reduce the channel dimension to finally generate the machine-vision-friendly frame $\hat{x}_{t}$. The above process is formulated as follows:
$f_{\textit{semantic}}=U(f_{\textit{MVS}},c_{1}),$ (8)
$\hat{x}_{t}=Conv2d(U(\alpha\cdot f_{\textit{semantic}}+(1-\alpha)\cdot f_{\textit{pixel}})),$
where $U(\cdot)$ denotes U-Net(s), $Conv2d(\cdot)$ denotes a 2D-convolutional layer, $\alpha$ denotes the extracted coefficient map for feature fusion, and $c_{1}$ denotes motion condition [29].
III-D Framework Optimization
TABLE I: Training strategy for proposed SEC-VCM framework. “M” stands for motion-related modules, “C” stands for contextual-coding-related modules, and “S” indicates proposed modules for machine-oriented reconstruction. Each epoch contains about 64k iterations.
Stage Modules $\mathcal{L}$ lrlr GOP Epochs
Base M $D_{\textit{MSE}}(x,x_{m})$ 1e-4 2 1
C $D_{\textit{MSE}}(x,\bar{x})$ 1e-4 2 1
C $D_{\textit{MSE}}(x,\bar{x})$ 1e-4 3 1
C $D_{\textit{MSE}}(x,\bar{x})$ 1e-4 6 1
M $D_{\textit{MSE}}(x,x_{m}),R_{m}$ 1e-4 6 8
C $D_{\textit{MSE}}(x,\bar{x}),R_{c}$ 1e-4 6 8
[…]
VCM S $\mathcal{L}_{e},\mathcal{L}_{\textit{cons}}$ 1e-4 2 1
S $\mathcal{L}_{e},\mathcal{L}_{\textit{cons}}$ 1e-4 3 1
S $\mathcal{L}_{e},\mathcal{L}_{\textit{cons}}$ 1e-4 6 1
[…]
M,C,S $\mathcal{L}_{\textit{total}}$ 1e-5 6 2
Using pixel-domain distortion metrics (e.g., MSE) rather than perception-related ones is essential in the early stage of codec training [4]. Meanwhile, existing optical flow-based inter-prediction methods are typically optimized for pixel-domain quality instead of MVS. Therefore, we divide the training process into two main stages. As shown in Table I, a multi-stage optimization strategy is adopted to achieve optimal rate–task performance.
In the base stage, the network is trained to achieve accurate inter-prediction and fine-grained pixel-domain reconstruction. This stage includes five sub-stages that separately optimize the motion-related and contextual-coding-related modules for improving overall rate-distortion (RD) performance, following [37, 29, 4]. To enhance perceptual quality, learned perceptual image patch similarity (LPIPS) [52] is also introduced, making the baseline codec more suitable for perception-driven tasks. After this stage, the codec is capable of completing the full compression loop. The rate-distortion loss $\mathcal{L}_{\textit{base}}$ of this stage is formulated as follows:
$\mathcal{L}_{\textit{base}}=\lambda_{\textit{MSE}}D_{\textit{MSE}}(x,\bar{x})+\lambda_{\textit{LPIPS}}D_{\textit{LPIPS}}(x,\bar{x})+R_{m}+R_{c},$ (9)
where $D_{\textit{MSE}}(\cdot,\cdot)$ denotes mean square error (MSE) distortion. $D_{\textit{LPIPS}}(\cdot,\cdot)$ denotes the LPIPS distortion. $R_{m}$ and $R_{c}$ denote bits per pixel (bpp) of the motion and context bitstreams, respectively. $\lambda_{\textit{MSE}}$ and $\lambda_{\textit{LPIPS}}$ denote weights for bitrate-distortion balance.
In the VCM stage, the SEC-VCM framework is optimized for machine vision with the help of visual backbones from two perspectives. First, the VFM is used to achieve hierarchical information symmetry from the entropy perspective (section III.B), formulated as:
$\mathcal{L}_{e}=H(f_{\textit{VB}}|f_{\textit{MVS}})+H(f_{\textit{MVS}}|f_{\textit{VB}}),$ (10)
where $H(\cdot|\cdot)$ denotes the conditional entropy. Second, multiple types of vision models are employed to maintain consistency between the reconstructed and original videos. Referring to [8, 6, 5], Spynet [35] is utilized to preserve pixel-level inter-frame relationships in reconstructed video, while DINO-v2 [22], Swin Transformer [53], and ResNet [21] are used to maintain machine-vision consistency in the feature domain. The consistency loss $\mathcal{L}_{\textit{cons}}$ is formulated as follows:
$),\mathcal{L}_{\textit{cons}}=\sum_{i}\sum_{j}\lambda_{i}\ D_{\textit{MSE}}(\textit{Model}_{i}^{j}(x),\textit{Model}_{i}^{j}(\hat{x}$)), (11)
where $\textit{Model}_{i}^{j}(\cdot)$ denotes the intermediate feature of the $j$-th layer in the $i$-th model, and $\lambda_{i}$ denotes the balance weight for the $i$-th model. Overall, the whole training strategy can be regarded as a multi-task optimization, as follows:
$\mathcal{L}_{\textit{total}}=\mathcal{L}_{\textit{base}}+\lambda_{e}\mathcal{L}_{e}+\lambda_{\textit{cons}}\mathcal{L}_{\textit{cons}},$ (12)
where $\lambda_{e}$ and $\lambda_{\textit{cons}}$ denote the balance weights.
IV Experiments
IV-A Experimental Settings of Video Codec
IV-A1 Settings of Vision Models
For the BiEC mechanism, Swin Transformer (base) [53] is employed during the process of training to align information with the proposed semantic decoder. Specifically, the intermediate features from its second, third, and fourth stages are utilized in our framework and named $f_{\textit{VB}}^{1/4}$, $f_{\textit{VB}}^{1/8}$, and $f_{\textit{VB}}^{1/16}$. To maintain machine-perceptual consistency, $res2$ and $res3$ from both ResNet-18 [21] and Swin Transformer (tiny) [53] are employed, along with the $4$-th and $8$-th layers of DINO-v2 (small) [22]. Deeper layers in VFM are excluded to prevent overfitting [8]. Additionally, the reconstructed frames are $2\times$ down-sampled by linear interpolation before being fed into Spynet [35], to prevent overemphasis on pixel-level information during training.
IV-A2 Training Datasets
The SEC-VCM is trained on the dataset merged from Vimeo-90k [54]. The Vimeo-90k dataset is widely used for training NVC frameworks and contains $89,800$ video sequences with 7 frames.
IV-A3 Training Settings
As shown in TABLE I, our training process consists of 50 epochs, with 40 epochs in the base stage and 10 epochs in the VCM stage. The training is conducted on four parallel NVIDIA RTX 3090 GPUs with a global batch size of 8, taking about 20 days. Notably, when training with long GOP, we retain gradients across continuous P-frames to better suppress the cumulative errors introduced by the lossy compression loop. Weight $\lambda_{\textit{MSE}}$ is randomly sampled from $[64,512]$. Since $D_{\textit{LPIPS}}$ and $\mathcal{L}_{\textit{cons}}$ are considered as distortion loss, their weights $\lambda_{\textit{LPIPS}}$ and $\lambda_{\textit{cons}}$ are set to the same value as $\lambda_{\textit{MSE}}$. The weight of the BiEC mechanism $\lambda_{e}$ is set to $1$.
IV-B Experimental Settings of Downstream Tasks
To validate the effectiveness of our SEC-VCM on downstream machine tasks, four types of downstream tasks with different levels of granularity are selected for evaluation.
IV-B1 Object Detection
Faster R-CNN [55] is implemented through the Detectron2 codebase [56] and tested on the YTVIS2019 dataset [57], which contains 2,883 videos in 40 categories, ranging in resolution from 360p to 1080p. Since Faster R-CNN has no officially released weights for the YTVIS2019 dataset. We fine-tune pretrained Faster R-CNN on the YTVIS2019 dataset using the default configuration [^1] Configuration of Faster R-CNN-R50: https://github.com/facebookresearch/detectron2/blob/main/configs/COCO-Detection/faster_rcnn_R_50_FPN_1x.yaml for the following experiments.
IV-B2 Video Instance Segmentation
The classical CNN-based model CrossVIS [^2] Configuration of CrossVIS-R50: https://github.com/hustvl/CrossVIS/blob/main/configs/CrossVIS/R_50_1x.yaml [58] and the Swin Transformer-based model Mask2Former-video [^3] Configuration of Mask2Former-Video-Swin-Base: https://github.com/facebookresearch/Mask2Former/blob/main/configs/youtubevis_2019/swin/video_maskformer2_swin_base_IN21k_384_bs16_8ep.yaml [59] are used to evaluate VIS performance. Notably, we directly adopt their officially released configurations and pretrained models. Experiments are conducted on the YTVIS2019 dataset.
IV-B3 Video Object Segmentation
We employ XMem [60] to evaluate the VOS performance on compressed videos. The DAVIS2017 dataset [61] is used in our experiments. It contains 150 videos with frame-level VOS annotations. The officially released 1080p videos are used for compression and subsequently resized to 480p for VOS.
IV-B4 Multiple Object Tracking
We employ officially released ByteTrack [62] on the MOT17 dataset [63], which contains 42 long videos with frame-level annotations and resolutions ranging from 480p to 1080p.
IV-C Comparison Methods
The proposed SEC-VCM is compared with four types of codecs. (1) Traditional hybrid codecs, including VTM-23.1 [1] and HM-18.0 [64]. Referring to VCM-related works [6, 8], these codecs are employed in the Low-Delay-P (LDP) mode. (2) Open-source neural video coding (NVC) frameworks, such as DCVC-RT [13], DCVC-FM [2], DCVC-DC [14], and DCVC-HEM [29]. (3) Human-perception-oriented and generative codecs I2VC [65], PLVC [40], and Control-GIC [66]. (4) Machine-oriented image codec SA-ICM [7] and video codec SMC++ [6]. The intra-period of PLVC is set to 9 following its official configuration, and the intra-period of other video codecs is set to 32 for fair comparison [29, 14]. All compared methods are implemented through their officially released weights.
IV-D Evaluation Metrics
BD-Rate [67] is used to evaluate the rate-task performance of the image and video codec. It measures the average bitrate difference required to achieve the same task performance relative to an anchor codec, where a lower value indicates better compression efficiency. VTM-23.1 serves as the anchor for calculating BD-Rate. In addition, bits per pixel (bpp) reflects the bitrate consumption of videos. A lower bpp value means a higher compression ratio. Task performance is evaluated by the metrics of each downstream task.
TABLE II: BD-rate (%) $\downarrow$ of compared video coding methods on five downstream machine vision task models. VTM-23.1 is the anchor for calculating BD-rate (%). All methods are split into four types: (1) standard codecs, (2) neural codecs for distortion in the pixel domain, (3) neural codecs for perception, and (4) neural codecs for machines. The best performance is displayed in red, and the runner-up performance is marked in blue.
Methods Types Faster R-CNN [55] CrossVIS [58] Mask2Former [59] Byte-track [62] XMem [60]
AP AP50 AP75 AP AP50 AP75 AP AP50 AP75 MOTA IDF IDs [^4] Experiments show that there is a weak correlation between the IDs metric and bitrate, as detailed in Fig. 12. Therefore, the IDs-based BD-rate does not well reflect the rate-task performance. J&F J F
VTM-23.1 [1] Standard 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
HM-18.0 [64] -2.2 1.3 -1.9 7.2 -15.5 -8.7 19.4 82.5 -21.3 11.3 32.7 -85.8 21.6 28.3 15.2
DCVC-HEM (ACM-MM’22) [29] 101.0 12.3 83.2 129.1 49.5 207.5 90.2 133.8 -3.9 -17.6 -29.7 -65.1 -2.6 -5.5 3.1
DCVC-DC (CVPR’23) [14] NVC 115.9 152.3 98.1 177.5 153.6 170.9 5.7 185.7 -15.9 5.7 -44.8 -78.4 4.6 -12.9 -6.0
DCVC-FM (CVPR’24) [2] (pixel) 108.2 137.2 90.9 127.1 25.2 186.8 33.4 110.9 -24.8 -23.5 -46.8 -95.5 44.3 48.6 39.4
DCVC-RT (CVPR’25) [13] 128.0 149.5 113.5 144.0 54.6 165.4 28.9 102.4 0.2 -23.0 -41.0 -94.5 87.1 89.5 85.8
PLVC (IJCAI’22) [40] NVC 130.9 143.3 117.0 63.4 20.9 -45.0 276.1 126.4 217.7 55.3 76.4 -72.0 362.5 402.7 277.3
I2VC (ArXiv’24) [65] (perception) 129.9 170.3 132.3 100.9 89.9 138.5 407.7 246.8 146.8 599.1 2292.9 -41.3 103.5 118.6 73.3
SA-ICM (ICIP’24) [8] [^5] The VCM-related image-codec SA-ICM is included in the comparison since open-sourced VCM methods are few, though this is kind of unfair. VCM 307.1 318.6 293.9 424.3 342.9 387.1 311.3 377.4 310.7 414.4 828.7 -182.3 160.4 165.8 156.7
SMC++ (ArXiv’24) [6] -13.1 -4.0 -20.6 -18.1 -39.0 -16.0 -32.2 -3.5 -0.1 91.8 -5.5 -72.4 -18.7 -21.7 -16.8
SEC-VCM (Ours) -46.2 -44.4 -47.6 -51.2 -64.8 -55.1 -23.6 62.9 -23.0 -44.9 17.9 -86.0 -29.8 -19.2 -38.4
Figure 7: Rate-task curve of object detection by Faster R-CNN [55] in terms of AP $\uparrow$ (1st column), AP50 $\uparrow$ (2nd column), and AP75 $\uparrow$ (3rd column). SEC-VCM achieves the best rate-task performance in all three metrics.
Figure 8: Visualization of detection results in YTVIS2019 dataset. Faster R-CNN [55] successfully detects the motion-blurred hand (upper row) and the small surfboard (lower row) from frames that are reconstructed by the proposed SEC-VCM.
Figure 9: Rate-task curve of video instance segmentation by CrossVIS [58] (1st row) and Mask2Former [59] (2nd row) in terms of AP $\uparrow$ (1st column), AP50 $\uparrow$ (2nd column), and AP75 $\uparrow$ (3rd column). SEC-VCM achieves the best rate-task performance on CrossVIS and secures the runner-up position in rate-AP and rate-AP75 performance on Mask2Former.
IV-E Rate-task Performance and Visualization
IV-E1 Object Detection
The average precision (AP) under three thresholds [55] is used to evaluate the performance of object detection. The proposed SEC-VCM achieves the best rate-task performance, as shown in TABLE II. It reduces $46.22\%$ (AP), $44.35\%$ (AP50), and $47.55\%$ (AP75) bitrate compared to VTM. The detection performance of all compared methods across various bitrates is shown in Fig. 7. The SEC-VCM consistently outperforms all methods in terms of detection AP, AP50, and AP75 at all tested bitrate points between $bpp=0.067$ and $bpp=0.211$.
We visualize the object detection results using Faster R-CNN [55], as shown in Fig. 8. On the reconstructed frames from SEC-VCM, Faster R-CNN successfully detects a motion-blurred hand and a small surfboard, which are missed in the original images and by several competing methods. This improvement is attributed to the fact that the perceptual-oriented reconstruction effectively removes artifacts and distortions that hinder detection.
Figure 10: Visualization on the YTVIS2019 dataset of video instance segmentation results. Under intense motion (1st row) and instance overlap (2nd row), Mask2Former [59] achieves more accurate instance contours on the reconstructed video from SEC-VCM.
IV-E2 Video Instance Segmentation
Figure 11: Rate-task curve of video object segmentation by XMem [60]. Task performance is evaluated in terms of J&F $\uparrow$ (1st column), J $\uparrow$ (2nd column), and F $\uparrow$ (3rd column). SEC-VCM has the best rate-J&F and rate-F performance and runner-up rate-J performance.
Figure 12: Rate-task curve of multiple object tracking by ByteTrack [62]. Task performance is evaluated in terms of MOTA $\uparrow$ (1st column), IDF1 $\uparrow$ (2nd column), and IDs $\downarrow$ (3rd column). SEC-VCM achieves the best rate-MOTA performance.
AP, AP50, and AP75 are evaluation metrics for VIS [58, 59]. As demonstrated in TABLE II and Fig. 9, when CrossVIS is used for instance segmentation, our proposed SEC-VCM achieves the highest compression ratio, saving $51.2\%$ (AP), $64.8\%$ (AP50), and $55.1\%$ (AP75) bitrate compared to VTM. When Mask2Former-video is used, SEC-VCM achieves the 2nd-best results in terms of AP and AP75, while SMC++ [6] achieves the best rate-task performance in terms of AP and AP50.
The visualizations of VIS by Mask2Former [59] are shown in Fig. 10. For the skateboard video with fast motion in the first row, Mask2Former preserves the object contours more accurately and suppresses incorrect segmentation on the video reconstructed by the proposed SEC-VCM. For the overlapping persons in the second row, it also provides more precise delineation on the reconstructed videos of SEC-VCM.
IV-E3 Multiple Object Tracking
Multiple object tracking accuracy (MOTA), ID F1 score (IDF1), and ID switches (IDs) are used to evaluate the performance of tracking, referring to [62]. As shown in TABLE II, our proposed SEC-VCM outperforms other methods in MOTA and achieves $44.9\%$ bitrate saving compared to VTM. Since IDF and IDs are mainly influenced by the inter-frame consistency rather than the frame quality, they show weaker correlation with bitrate. Although SEC-VCM does not have the best rate-task performance in terms of IDF1 and IDs, it outperforms other methods on variable bitrates, as shown in Fig. 12.
IV-E4 Video Object Segmentation
Three metrics—Jaccard value (J), F-measure (F), and J&F—are used to evaluate the performance of video object segmentation. As demonstrated in TABLE II and Fig. 11, our proposed SEC-VCM achieves the most bitrate saving compared to VTM in terms of J&F ($29.83\%$) and F ($38.42\%$) while achieving a comparable compression ratio with SMC++ on the metric of J.
IV-E5 Discussion
The above experiments show that the SEC-VCM achieves promising rate-task performance in most metrics. It is worth mentioning that although many methods exhibit excellent rate-distortion (PSNR, MS-SSIM) and rate-perception (LPIPS, etc.) performance that far exceeds standard codecs, their performance in downstream machine vision tasks is not satisfactory. These widely used metrics cannot well reflect the needs of machine vision tasks [68], which further illustrates the necessity of VCM-related works.
TABLE III: Influence of different backbones on BiEC mechanism. VTM-23.1 is the anchor for calculating BD-rate (%) $\downarrow$. The best rate-task performance is displayed in red.
Downstream tasks Metric SwinTransformer ResNet VGGNet
Faster R-CNN [55] AP -46.2 -49.0 -46.3
CrossVIS [58] AP -51.2 -32.8 -30.0
Mask2Former [59] AP -23.6 -18.9 -35.4
XMem [60] J&F -29.8 3.8 0.1
ByteTrack [62] MOTA -44.9 -23.7 -20.9
Average BD-Rate -39.2 -24.1 -26.5
IV-F Ablation Study
IV-F1 Ablation on the Guidance Visual Backbone in BiEC Mechanism
While many commonly used backbones have demonstrated strong performance across a wide range of computer vision tasks, we integrate different types of backbones into the proposed BiEC mechanism. Since video reconstruction is a hierarchical up-sampling process, three backbones with hierarchical feature extraction capabilities—Swin Transformer [53], ResNet [21], and VGGNet [69]—are employed in the BiEC mechanism. The evaluation is conducted on the primary metrics of five downstream tasks. As shown in Table III, the results reveal notable performance differences across frameworks. The Swin Transformer is eventually chosen as the guidance backbone in the BiEC mechanism, as it significantly outperforms ResNet and VGGNet in three downstream tasks.
TABLE IV: Influence of the BiEC loss weight $\lambda_{e}$. $H$ denotes the value of conditional entropy. VTM-23.1 is the anchor for calculating BD-rate (%) $\downarrow$. The best rate-task performance is displayed in red.
Downstream tasks Metric $\lambda_{e}$=$0.1$ $\lambda_{e}$=$0.5$ $\lambda_{e}$=11 $\lambda_{e}$=55 $\lambda_{e}$=$10$
HH=$1.73$ HH=$1.69$ HH=$1.59$ HH=$1.50$ HH=$1.45$
Faster R-CNN [55] AP -40.2 -45.8 -46.2 -41.0 -44.4
CrossVIS [58] AP -31.5 -54.4 -51.2 -38.8 -46.5
Mask2Former [59] AP -33.3 -11.6 -23.6 -33.9 -27.2
XMem [60] J&F -16.9 1.4 -19.2 -38.4 -25.9
ByteTrack [62] MOTA -10.4 -29.5 -44.9 -16.5 -9.7
Average BD-Rate -24.5 -26.7 -37.0 -33.7 30.8
TABLE V: Ablation study on the BiEC-related components.
Models $H(f_{\textit{MVS}}|f_{\textit{VB}})$ $H(f_{\textit{VB}}|f_{\textit{MVS}})$ Multi-layer Average bitrate
Ours 0%
$M_{1}$ +16.0%
$M_{2}$ +7.4%
$M_{3}$ +11.3%
$M_{4}$ +26.9%
$M_{5}$ +21.5%
$M_{6}$ MSE MSE +21.1%
$M_{7}$ KL KL +17.3%
IV-F2 Ablation on the BiEC Mechanism
Although visual backbones exhibit generalization ability, it is crucial to prevent the entire codec from overfitting to a specific backbone. To evaluate the impact of different levels of the BiEC mechanism on downstream task performance, we train the SEC-VCM with different $\lambda_{e}$, as shown in Table IV. Finally, $\lambda_{e}$ is set to achieve a good trade-off across multiple downstream tasks.
We conduct ablation studies on the bi-directional design and the multi-layer design, as shown in Table V. Results from $M_{1}$ and $M_{2}$ demonstrate that each direction of the entropy constraints plays a critical role in improving rate-task performance (average 11.7%). Also, the experimental results of $M_{3}$, $M_{4}$, and $M_{5}$ show that removing multi-layer design results in significant rate-task performance degradation (average 12.0%). Notably, in $M_{6}$, the proposed conditional entropy loss $\mathcal{L}_{e}$ is replaced with MSE, leading to a 21.1% decrease in rate-task performance, highlighting the superiority of soft alignment over rigid matching.
TABLE VI: Ablation study on the semantic-pixel dual-path fusion.
Models Gated fusion Concatenation Addition Average bitrate
Ours 0%
$M_{8}$ +13.4%
$M_{9}$ +8.9%
$M_{10}$ +19.5%
TABLE VII: Complexity of neural codecs in terms of inference time, encoding time, decoding time, and number of optimized parameters. The times are evaluated on 1080p videos by a single NVIDIA RTX 3090.
Methods Inference Encoding Decoding Optimized
(ms) (ms) (ms) params
DCVC-HEM (ACM-MM’22) [29] 566 582 209 17.5M
DCVC-DC (CVPR’23) [14] 670 701 528 19.8M
DCVC-FM (CVPR’24) [2] 654 684 537 18.3M
DCVC-RT (CVPR’25) [13] [^6] We turn off its self-implemented CUDA operations and only enable PyTorch-based operations for fair comparison. 81 36 39 20.7M
PLVC (IJCAI’22) [40] [^7] Since the officially released code of PLVC is not supported by the NVIDIA RTX 3090 and the corresponding CUDA version, it is tested on CPU. The discriminator from its GAN architecture is not included in the number of parameters. / / / 19.2M
I2VC (ArXiv’24) [65] 1030 / / 78.4M
Control-GIC (ICLR’25) [66] / 847 1285 131.0M
SA-ICM (ICIP’24) [7] 720 647 768 76.5M
SMC++ (ArXiv’24) [6] 1188 / / 96.2M
SEC-VCM (Ours) 707 584 411 31.9M
Figure 13: Visualization of the intermediate results in the gated fusion process. The pixel-oriented feature $f_{\textit{pixel}}$ and machine-oriented feature $f_{\textit{semantic}}$ are selectively combined based on $\alpha$ for the final reconstruction. Red and blue indicate high and low feature values, respectively.
IV-F3 Ablation on the SPDF Module
To explore how to effectively utilize pixel-oriented features $f_{\textit{pixel}}$ for MVS-oriented reconstruction, channel-wise concatenation ($M_{8}$) and direct addition ($M_{9}$) are employed to replace the gated fusion, as shown in Table VI. Experimental results demonstrate that gated fusion outperforms the other two fusion methods. Notably, the rate-task performance drops significantly by $19.5\%$ when $f_{\textit{pixel}}$ is not used ($M_{10}$). Furthermore, visualization results indicate that the gated fusion mechanism can selectively combine $f_{\textit{semantic}}$ and $f_{\textit{pixel}}$ for final reconstruction. As shown in Fig. 13, $f_{\textit{semantic}}$ plays the major role in the reconstruction on channel 9, while contributing only features related to boats, forests, lake surface, and sky on channels 21, 25, and 36.
IV-G Complexity Analysis
We compare the inference time, encoding time, decoding time, and the number of optimized parameters of the proposed SEC-VCM with other neural codecs [29, 14, 2, 65, 66, 6, 7], as shown in Table VII. The inference time reflects the overall computational complexity of the model, excluding file I/O. The encoding and decoding times represent the complexity at the encoder and decoder sides, respectively. The inference, encoding, and decoding times are recorded on 1080p videos and a single NVIDIA RTX 3090 with a batch size of 1. In terms of inference, encoding, and decoding times, our method is comparable to existing high-performance neural video codecs [29, 14, 2] and significantly faster than existing human-perception-oriented codecs [65, 66] and machine-vision-oriented codecs [7, 6]. In terms of the number of optimized parameters, our framework is lighter than machine-vision-oriented codecs SA-ICM [7] and SMC++ [6]. Still, it is larger than other neural video codecs due to additional modules (semantic decoder and SPDF module) for machine-oriented reconstruction.
V Conclusion
In this paper, we propose a unified Symmetric Entropy-Constrained Video Coding framework for Machines (SEC-VCM). By aligning the video coding process with the video understanding process of the visual backbone from an entropy perspective, SEC-VCM eliminates machine-vision-irrelevant information. The proposed BiEC mechanism softly maintains consistency between the visual backbone’s encoding process and the video decoding process, while the semantic-pixel dual-path fusion module introduces pixel-level priors to suppress artifacts and enhance reconstruction quality. Extensive experiments on diverse downstream tasks demonstrate the superiority and generalization ability of our method in achieving MVS-friendly reconstruction. We hope our proposed method offers insights into bridging the gap between neural video coding and machine vision, serving as a step forward for advancing the field of video coding for machines (VCM).

### v2

@@
 As video transmission increasingly serves machine vision systems (MVS) instead of human vision systems (HVS), video coding for machines (VCM) has become a critical research topic.
-Existing VCM methods often bind codecs to specific downstream models, requiring retraining or supervised data and thus limiting generalization in multi-task scenarios.
+Existing VCM methods often bind codecs to specific downstream models, requiring retraining or supervised data, thus limiting generalization in multi-task scenarios.
 Recently, unified VCM frameworks have employed visual backbones (VB) and visual foundation models (VFM) to support multiple video understanding tasks with a single codec.
@@
 Specifically, a bi-directional entropy-constraint (BiEC) mechanism ensures symmetry between the process of video decoding and VB encoding by suppressing conditional entropy.
-This helps the codec to explicitly handle semantic information beneficial for MVS while squeezing useless information.
-Furthermore, a semantic–pixel dual-path fusion (SPDF) module injects pixel-level priors into the final reconstruction.
+This helps the codec to explicitly handle semantic information beneficial to MVS while squeezing useless information.
+Furthermore, a semantic-pixel dual-path fusion (SPDF) module injects pixel-level priors into the final reconstruction.
 Through semantic-pixel fusion, it suppresses artifacts harmful to MVS and improves machine-oriented reconstruction quality.
-Experimental results show our framework achieves state-of-the-art (SOTA) in rate–task performance, with significant bitrate savings over VTM on video instance segmentation (37.41%), video object segmentation (29.83%), object detection (46.22%), and multiple object tracking (44.94%).
-We will release our code.
+Experimental results show our framework achieves state-of-the-art (SOTA) in rate-task performance, with significant bitrate savings over VTM on video instance segmentation (37.4%), video object segmentation (29.8%), object detection (46.2%), and multiple object tracking (44.9%).
+We will release our code soon.
 I Introduction
@@
 $I(\cdot)$ denotes mutual information, $H(\cdot)$ denotes Shannon entropy, and $\Uparrow$” denotes that the optimization goal is to maximize the value.
-$\mathcal{E}_{\textit{visual}}$ denotes the encoder of visual backbone (VB) or visual foundation model (VFM).
+$\mathcal{L}_{a}$ and $\mathcal{L}_{b}$ denote downstream task losses, $\mathcal{L}_{\textit{cons}}$ denotes loss for maintaining feature consistency, and $\mathcal{E}_{\textit{visual}}$ denotes the encoder of visual backbone (VB) or visual foundation model (VFM).
 (a) Standard or neural video codecs [1, 2] are optimized for HVS, which makes the decoded videos $\hat{x}$ suboptimal for machine vision.
-(b) Each neural video/feature codec [3, 4] is optimized for a specific machine vision task.
+(b) Each neural video/feature codec [3, 4, 5] is optimized for a specific machine vision task by downstream task loss.
 (c) The unified codec [6, 7, 8, 9] is optimized with the assistance of the VB or VFM to reconstruct videos suitable for multiple downstream tasks.
@@
 Meanwhile, videos collected for machine vision tasks, such as surveillance analysis [10] and facial recognition [11], typically require only part of the video information, rather than the complete visual details [12, 13].
-Currently, neural video coding frameworks have evolved significantly and offer excellent compression performance [13, 2, 14, 15, 16].
-However, these methods are optimized primarily for HVS-related metrics and do not fully satisfy the requirements of machine vision systems (MVS) in terms of video understanding (e.g., detection, segmentation, and tracking) [17], as shown in Fig.
+Currently, neural video coding frameworks have evolved significantly and offer excellent compression performance [14, 15, 16, 17, 18, 19].
+However, these methods are optimized primarily for pixel-domain distortion or HVS-related metrics and do not fully satisfy the requirements of machine vision systems (MVS) in terms of video understanding (e.g., detection, segmentation, and object tracking) [20], as shown in Fig.
 1(a).
@@
 Meanwhile, some parameter-efficient fine-tuning (PEFT) strategies are introduced to adapt codecs to new downstream tasks with lower cost [3].
-Nonetheless, most of them still rely on the labeled data and supervised training process for specific downstream tasks, limiting their generalization in multi-task scenario.
+Nonetheless, most of them still rely on the labeled data and supervised training process for specific downstream tasks, limiting their generalization in multi-task scenarios.
 To improve the applicability, some methods [23, 9] have adopted unified compression frameworks to reconstruct videos suitable for multiple downstream tasks simultaneously.
@@
 1(d).
-Its key principle is to bridge the gap between the video coding and understanding processes from an entropy perspective, explicitly guiding the video codec to handle only information beneficial for video understanding while discarding irrelevant information.
+Its key principle is to bridge the gap between the video coding and understanding processes from an entropy perspective, explicitly guiding the video codec to handle only information beneficial to video understanding while discarding irrelevant information.
 Specifically, an information-alignment-based semantic decoder with a bi-directional entropy-constraint (BiEC) mechanism is proposed to make the video decoding process of the codec symmetric and consistent with the encoding process of the visual backbone.
@@
 The experiments on four machine vision tasks demonstrate that the proposed SEC-VCM achieves state-of-the-art (SOTA) rate-task performance.
-Compared to VTM, our method achieves significant bitrate reductions, with savings of 37.41% for video instance segmentation, 29.83% for video object segmentation, 46.22% for object detection, and 44.94% for multiple object tracking.
+Compared to VTM, our method achieves significant bitrate reductions, with savings of 37.4% for video instance segmentation, 29.8% for video object segmentation, 46.2% for object detection, and 44.9% for multiple object tracking.
 Our main contributions are summarized as follows:
@@
 II-A Neural Video Coding
-Most existing neural video coding methods [27, 28, 29, 14, 30, 31, 32, 15, 16, 33, 34] follow the motion-then-residual paradigm and mainly include inter-prediction and residual/context compression.
+Most existing neural video coding methods [29, 30, 31, 16, 32, 17, 18, 33] adopt a two-step pipeline including inter-prediction and residual or context compression.
 Lu et al.
-[27] proposed the first end-to-end video coding framework DVC, which uses the optical flow [35] for inter-prediction.
-Agustsson et al.
-[30] proposed the Scale Space Flow and added a Gaussian dimension to create more reference pixels for inter-frame prediction.
+[29] proposed the first neural video codec DVC, which uses the optical flow [34] for inter-prediction.
 Hu et al.
-[36] proposed FVC to convert the video from the pixel domain to the feature domain and used deformable convolution for ME-MC in the feature domain.
+[35] proposed to convert the processes of motion estimation (ME), motion compensation (MC), and residual coding from the pixel domain to the feature domain.
 Li et al.
-[28] improved the concept of predictive coding and proposed the conditional coding paradigm for better compression performance, termed DCVC.
+[30] improved the concept of residual coding and proposed the conditional coding paradigm for better compression performance, termed DCVC.
 Sheng et al.
@@
 Furthermore, many methods [17, 15, 19] introduce long-range training to suppress accumulation error and make the codec perform better on a long group of pictures (GOP).
-Overall, existing neural video coding methods improve compression efficiency from various perspectives, such as inter- or bi-directional prediction [30, 15, 14, 16, 39, 40], residual/context compression [13], and entropy models [29, 13].
-Many NVC methods, such as ECVC-series [38] and DCVC-series [14, 2, 13, 15, 39], demonstrate excellent compression performance.
+Overall, existing neural video coding methods improve compression efficiency from various perspectives, such as inter- or bi-directional prediction [32, 17, 16, 18, 37], residual/context compression [14], and entropy models [31, 14].
+Many NVC methods, such as ECVC-series [19] and DCVC-series [16, 15, 14], demonstrate excellent compression performance.
 II-B Neural Video Coding for Machines
@@
 (2) The semantic decoding process is information-aligned with the encoding process of the pretrained visual backbone by the proposed BiEC mechanism from an entropy perspective.
-This helps the semantic decoder $\mathcal{D}_{\textit{sem}}$ purely handle information beneficial for MVS.
+This helps the semantic decoder $\mathcal{D}_{\textit{sem}}$ purely handle information beneficial to MVS.
 (3) The pixel-level prior knowledge $f_{\textit{pixel}}$ is injected into the perception-oriented decoding process to fulfill frame content, suppress artifacts, and enhance final reconstruction quality.
@@
 Current methods optimize the codec for specific or multiple machine vision tasks, either in a supervised or unsupervised manner [4, 5, 6, 7].
-Also, some methods maintain semantic consistency by minimizing the perceptual distance between the original video $x$ and compressed video $\hat{x}$ measured by visual backbone (VB) or visual foundation model (VFM) [8, 7].
+Also, some methods maintain semantic consistency by minimizing the perceptual distance between the original video $x$ and the compressed video $\hat{x}$ measured by visual backbone (VB) or visual foundation model (VFM) [9, 8].
 Their paradigm is formulated as follows:
 […]
-where $\mathcal{E}(\cdot)$ and $\mathcal{D}(\cdot)$ denote the encoder and decoder of the video codec, $\lfloor\cdot\rceil$” denotes the quantization operation.
-$\mathcal{L}$ denotes the total loss, and $R$ denotes the bitrate.
-$\mathcal{E}_{\textit{visual}}$ denotes the encoder of VB or VFM.
-$D(\cdot,\cdot)$ denotes pixel-domain or feature-domain distortion (e.g., MSE).
-Although VB and VFM have been introduced to help codecs preserve semantics, their integration with video coding remains insufficiently explored.
+where $\mathcal{E}(\cdot)$ and $\mathcal{D}(\cdot)$ denote the encoder and decoder of the video codec, $\lfloor\cdot\rceil$” denotes the quantization operation [46].
+$\mathcal{L}$ denotes the total loss, $R$ denotes the bitrate, $\mathcal{E}_{\textit{visual}}$ denotes the encoder of VB or VFM, $D(\cdot,\cdot)$ denotes pixel-domain or feature-domain distortion (e.g., MSE), and $\mathcal{L}_{task(\hat{x})}$ denotes the loss function of the downstream task.
+Although VB and VFM have been introduced to help codecs preserve semantics, their integration with video coding remains unexplored.
 They excel at extracting rich and general feature representations for multiple machine vision tasks [24, 25], yet such features are not naturally suited for entropy coding or direct compression.
-To better leverage their general capacity, we align the video codec with VB/VFM from an entropy perspective, encouraging it to selectively preserve features beneficial for MVS while squeezing irrelevant information.
+To better leverage their general capacity, we propose to align the video codec with VB/VFM from an entropy perspective, encouraging it to selectively preserve features beneficial to MVS while squeezing irrelevant information.
 The alignment objective is formulated as:
 $\mathcal{D}_{\textit{sem}}(\hat{y})\xleftrightarrow{\text{aligned}}\mathcal{E}_{\textit{visual}}(x)\xrightarrow{\text{capable}}\text{downstream tasks},$ (2)
-where $\mathcal{D}_{\textit{sem}}(\cdot)$ denotes the semantic decoder and $\hat{y}$ is the transmitted bitstream (entropy-coded latent representation).
+where $\mathcal{D}_{\textit{sem}}(\cdot)$ denotes the semantic decoder and $\hat{y}$ denotes the transmitted bitstream (entropy-coded latent representation).
 Additionally, completing the compression loop requires low-level details to support inter-prediction [4].
 Therefore, we propose that pixel-oriented reconstruction and machine-oriented reconstruction are conducted separately, with the former providing pixel-level priors for the latter.
-Figure 3: Structures of proposed modules.
-(a) Structure of the information-alignment-based semantic decoder and BiEC mechanism.
+Figure 3: (a) Structure of the proposed information-alignment-based semantic decoder and BiEC mechanism.
 DEDE denotes the distribution estimation module, and $\uparrow\times 2$” denotes two times up-sampling.
 During training, the intermediate features for the decoding $f_{\textit{MVS}}$ are made hierarchically symmetric with the features $f_{\textit{VB}}$ from the visual backbone.
-(b) Structure of the SPDF module.
+(b) Structure of the proposed SPDF module.
 Machine vision system (MVS)-oriented semantic feature $f_{t}^{\textit{MVS}}$ and pixel-related feature $f_{t}^{\textit{pixel}}$ are selectively combined based on the factor $\alpha$.
@@
 Specifically, it consists of three main steps: inter-prediction and contextual coding for the basic compression loop, machine-oriented semantic decoding with the Bi-directional Entropy-Constraint (BiEC) mechanism, and semantic-pixel gated fusion.
-(1) In the first step, we inherit the basic compression loop from DCVC-HEM [29], where motion and contextual information are compressed, quantized, entropy-coded, and then reconstructed.
+(1) In the first step, we adopt the basic compression loop from DCVC-HEM [31], where motion and contextual information are compressed, quantized, entropy-coded, and then reconstructed.
 The compensated features $c_{1}$, $c_{2}$, and $c_{3}$ serve as motion priors for both contextual coding and our semantic decoding.
@@
 Specifically, in the BiEC mechanism, the symmetry is enforced at three scales, and the alignment is achieved by constraining conditional entropy.
-(3) In the third step, since $f_{\textit{MVS}}$ mainly contains the semantic information, additional detailed pixel-level priors $f_{\textit{pixel}}$ are injected into the final reconstruction process to enrich frame details, suppress artifacts, and produce frames better suited for MVS.
-Notably, our framework is general across multiple downstream tasks, and we therefore remove the task loss $\mathcal{L}_{task(\hat{x})}$ that relies on specific labeled datasets.
+(3) In the third step, since $f_{\textit{MVS}}$ mainly contains the semantic information, additional detailed pixel-level priors $f_{\textit{pixel}}$ are injected into the final reconstruction process to enrich frame details, suppress artifacts, and produce frames better suited to MVS.
+Notably, our framework is general across multiple downstream tasks; therefore, we remove the task loss $\mathcal{L}_{task(\hat{x})}$ that relies on specific labeled datasets.
 Our paradigm is illustrated as follows:
@@
 (3)
-where $\mathcal{D}_{\textit{sem}}(\cdot)$ denotes the proposed information-alignment-based semantic decoder, $f_{\textit{VB}}$ denotes the feature extracted by the visual backbone, $f_{\textit{MVS}}$ denotes the feature in the semantic decoder, and $f_{\textit{pixel}}$ denotes the feature for low-level reconstruction.
+where $\mathcal{D}_{\textit{sem}}(\cdot)$ denotes the proposed information-alignment-based semantic decoder, $f_{\textit{VB}}$ denotes the feature extracted by the visual backbone, $f_{\textit{MVS}}$ denotes the feature in the semantic decoder, and $f_{\textit{pixel}}$ denotes the feature used for low-level reconstruction.
 $H(f_{\textit{MVS}}|f_{\textit{VB}})$ and $H(f_{\textit{VB}}|f_{\textit{MVS}})$ denote the conditional entropy terms used as loss items in the proposed BiEC mechanism.
 III-B Semantic Decoding with Symmetric Entropy-Constraint
-To leverage the visual backbone’s strong capability of extracting machine-vision-oriented feature representations, the video decoding process is hierarchically and symmetrically aligned with the encoding process of VB during training.
+To leverage the visual backbone’s strong capability for extracting machine-vision-oriented feature representations, the video decoding process is hierarchically and symmetrically aligned with the encoding process of VB during training.
 It enables the decoder to inherit the prior knowledge learned by the backbone that has experienced large-scale pretraining, focusing purely on features relevant to machine vision tasks.
@@
 It progressively recovers the quantized and compact latent representation $\hat{y}_{ctx}$ (at $1/16$ resolution) back to the original spatial resolution, producing machine-vision-oriented features $f_{\textit{MVS}}$.
-The inter-predicted features $c_{2}$ ($1/2$ size) and $c_{3}$ ($1/4$ size) are used as conditions for frame reconstruction, referring to [28, 29].
+The inter-predicted features $c_{2}$ (at $1/2$ resolution) and $c_{3}$ (at $1/4$ resolution) are used as conditions for frame reconstruction, referring to [30, 31].
 Notably, intermediate features $f_{\textit{MVS}}$ within the semantic decoder are aligned with the intermediate results $f_{\textit{VB}}$ from the visual backbone at three distinct scales.
@@
 $\Downarrow$” indicates that the optimization goal is to minimize the item.
-Figure 5: The t-SNE visualization of intermediate features $f_{\textit{MVS}}^{1/8}$ in semantic decoder with and without symmetric information alignment.
+Figure 5: The t-SNE visualization of intermediate features $f_{\textit{MVS}}^{1/8}$ in the semantic decoder without and with symmetric information alignment.
 After alignment via the BiEC mechanism, intra-class distances are effectively reduced, and the intermediate features $f_{\textit{MVS}}^{1/8}$ in the decoding process exhibit more discriminative semantic representations.
-To enforce symmetry, we further impose a conditional entropy constraint in the reverse direction, making $f_{\textit{VB}}$ a subset of $f_{\textit{MVS}}$.
+To enforce symmetry, we further impose a conditional entropy constraint in the reverse direction, thus making $f_{\textit{VB}}$ a subset of $f_{\textit{MVS}}$.
 Through this bi-directional constraint, the features $f_{\textit{MVS}}$ in the semantic decoder and the visual-backbone-extracted representations $f_{\textit{VB}}$ are softly aligned, as illustrated in Fig.
@@
 where $\mu_{\textit{VB}}^{i}$ and $\sigma_{\textit{VB}}^{i}$ denote estimated element-wise distribution parameters of $f_{\textit{VB}}^{i}$.
-Specifically, each distribution estimation module contains a convolution layer and a ResBlock [21].
+Specifically, each DE module contains a convolutional layer and a ResBlock [24].
 Estimated distribution parameters $\mu$ and $\sigma$ have the same shape as the corresponding feature $f$.
@@
 Figure 6: Visualization results of reconstructed frame using different feature types at $bpp\approx 0.08$.
-Relying only on pixel-oriented feature $f_{\textit{pixel}}$ tends to miss fine details of the dog’s fur textures, while using only semantic-level feature $f_{\textit{semantic}}$ leads to the loss of grass texture and introduces artifacts.
+Relying only on pixel-oriented feature $f_{\textit{pixel}}$ tends to miss fine details of the dog’s fur texture, while using only semantic-level feature $f_{\textit{semantic}}$ leads to the loss of grass texture and introduces artifacts.
 By fusing $f_{\textit{semantic}}$ and $f_{\textit{pixel}}$, our method reconstructs the dog’s fur details better and avoids artifacts in the grass.
 III-B3 Discussion
-Through the BiEC mechanism, the semantic decoder is encouraged to reconstruct semantic representations by relying only on features that are beneficial for video understanding, guided by the visual backbone.
+Through the BiEC mechanism, the semantic decoder is encouraged to reconstruct semantic representations by relying only on features that are beneficial to video understanding, guided by the visual backbone.
 In other words, the intermediate features $f_{\textit{MVS}}$ of the semantic decoder are expected to be semantic-aware.
@@
 Specifically, 400 images from 8 classes of the ImageNet-1k [24] validation set are fed into our framework, and the intermediate features $f_{\textit{MVS}}^{1/8}$ of the semantic decoder are projected into a two-dimensional space using t-SNE [47].
-Since our framework is designed for P-frames and cannot be evaluated on a single frame, the reference frame is set to all zeros for this visualization.
+Since our framework is designed for P-frames and cannot be directly evaluated on single frames, all pixels of the reference frame are set to zero for this visualization.
 As shown in Fig.
@@
 Specifically, as illustrated in Fig.
-6, using only $f_{\textit{MVS}}$ helps retain the texture of the dog’s fur and may increase its detection accuracy, but it ignores the texture of the grass and introduces visible artifacts.
-In contrast, relying solely on $f_{\textit{pixel}}$ leads to blur under limited bitrates.
-By gated fusing both $f_{\textit{MVS}}$ and $f_{\textit{pixel}}$, our method effectively suppresses grass artifacts and preserves fine textures of fur.
+6, using only $f_{\textit{MVS}}$ helps retain the texture of the dog’s fur and may increase its detection accuracy, but it ignores grass textures and introduces artifacts, which undermine the semantics of the grass region.
+In contrast, relying solely on $f_{\textit{pixel}}$ leads to a blur under limited bitrates.
+By fusing both $f_{\textit{MVS}}$ and $f_{\textit{pixel}}$, our method effectively suppresses grass artifacts and preserves fine textures of fur.
 The structure of the SPDF module is shown in Fig.
 3(b).
-Specifically, the machine-oriented feature $f_{\textit{MVS}}$ and pixel-level features $f_{\textit{pixel}}$ are both fed into the SPDF module.
-A residual block is employed to calculate an element-wise gating coefficient $\alpha$, which is formulated as follows:
+Specifically, the machine-oriented feature $f_{\textit{MVS}}$ and pixel-level feature $f_{\textit{pixel}}$ are both fed into the SPDF module.
+A residual block is employed to compute an element-wise gating coefficient $\alpha$, which is formulated as follows:
 $\alpha$ $=Sigmoid(ResBlock(Concat(f_{\textit{MVS}},f_{\textit{pixel}}))),$ (7)
@@
 $\hat{x}_{t}=Conv2d(U(\alpha\cdot f_{\textit{semantic}}+(1-\alpha)\cdot f_{\textit{pixel}})),$
-where $U(\cdot)$ denotes U-Net(s), $Conv2d(\cdot)$ denotes a 2D-convolutional layer, $\alpha$ denotes the extracted coefficient map for feature fusion, and $c_{1}$ denotes motion condition [29].
+where $U(\cdot)$ denotes U-Net(s), $Conv2d(\cdot)$ denotes a 2D-convolutional layer, $\alpha$ denotes the extracted coefficient map for feature fusion, and $c_{1}$ denotes the motion condition [31].
 III-D Framework Optimization
@@
 M,C,S $\mathcal{L}_{\textit{total}}$ 1e-5 6 2
-Using pixel-domain distortion metrics (e.g., MSE) rather than perception-related ones is essential in the early stage of codec training [4].
+Using pixel-domain distortion metrics (e.g., MSE) rather than perception-related metrics is essential in the early stage of codec training [4].
 Meanwhile, existing optical flow-based inter-prediction methods are typically optimized for pixel-domain quality instead of MVS.
 Therefore, we divide the training process into two main stages.
-As shown in Table I, a multi-stage optimization strategy is adopted to achieve optimal rate–task performance.
+As shown in Table I, a multi-stage optimization strategy is adopted to achieve optimal rate-task performance.
 In the base stage, the network is trained to achieve accurate inter-prediction and fine-grained pixel-domain reconstruction.
-This stage includes five sub-stages that separately optimize the motion-related and contextual-coding-related modules for improving overall rate-distortion (RD) performance, following [37, 29, 4].
+This stage includes five sub-stages that separately optimize the motion-related and contextual-coding-related modules to improve overall rate-distortion (RD) performance, following [36, 31, 4].
 To enhance perceptual quality, learned perceptual image patch similarity (LPIPS) [48] is also introduced, making the baseline codec more suitable for perception-driven tasks.
@@
 Deeper layers in VFM are excluded to prevent overfitting [9].
-Additionally, the reconstructed frames are $2\times$ down-sampled by linear interpolation before being fed into Spynet [35], to prevent overemphasis on pixel-level information during training.
+Additionally, the reconstructed frames are twice down-sampled by linear interpolation before being fed into Spynet [34], to prevent overemphasis on pixel-level information during training.
 IV-A2 Training Datasets
@@
 Weight $\lambda_{\textit{MSE}}$ is randomly sampled from $[64,512]$.
-Since $D_{\textit{LPIPS}}$ and $\mathcal{L}_{\textit{cons}}$ are considered as distortion loss, their weights $\lambda_{\textit{LPIPS}}$ and $\lambda_{\textit{cons}}$ are set to the same value as $\lambda_{\textit{MSE}}$.
+Since $D_{\textit{LPIPS}}$ and $\mathcal{L}_{\textit{cons}}$ are considered distortion losses, their weights $\lambda_{\textit{LPIPS}}$ and $\lambda_{\textit{cons}}$ are set to the same value as $\lambda_{\textit{MSE}}$.
 The weight of the BiEC mechanism $\lambda_{e}$ is set to $1$.
@@
 Faster R-CNN [51] is implemented through the Detectron2 codebase [52] and tested on the YTVIS2019 dataset [53], which contains 2,883 videos in 40 categories, ranging in resolution from 360p to 1080p.
-Since Faster R-CNN has no officially released weights for the YTVIS2019 dataset.
+Since Faster R-CNN has no officially released weights on the YTVIS2019 dataset.
 We fine-tune pretrained Faster R-CNN on the YTVIS2019 dataset using the default configuration [^1] Configuration of Faster R-CNN-R50: https://github.com/facebookresearch/detectron2/blob/main/configs/COCO-Detection/faster_rcnn_R_50_FPN_1x.yaml for the following experiments.
@@
 IV-B4 Multiple Object Tracking
-We employ officially released ByteTrack [62] on the MOT17 dataset [63], which contains 42 long videos with frame-level annotations and resolutions ranging from 480p to 1080p.
+We employ officially released ByteTrack on the MOT17 dataset [58], which contains 42 long videos with frame-level annotations and resolutions ranging from 480p to 1080p.
 IV-C Comparison Methods
@@
 IV-D Evaluation Metrics
-BD-Rate [67] is used to evaluate the rate-task performance of the image and video codec.
+BD-Rate [63] is used to evaluate the rate-task performance of the image and video codecs.
 It measures the average bitrate difference required to achieve the same task performance relative to an anchor codec, where a lower value indicates better compression efficiency.
@@
 Methods Types Faster R-CNN [51] CrossVIS [54] Mask2Former [55] Byte-track [58] XMem [56]
-AP AP50 AP75 AP AP50 AP75 AP AP50 AP75 MOTA IDF IDs [^4] Experiments show that there is a weak correlation between the IDs metric and bitrate, as detailed in Fig.
+AP AP50 AP75 AP AP50 AP75 AP AP50 AP75 MOTA IDF1 IDs [^4] Experiments show that there is a weak correlation between the IDs metric and bitrate, as detailed in Fig.
 12.
@@
 VCM 307.1 318.6 293.9 424.3 342.9 387.1 311.3 377.4 310.7 414.4 828.7 -182.3 160.4 165.8 156.7
-SMC++ (ArXiv’24) [6] -13.1 -4.0 -20.6 -18.1 -39.0 -16.0 -32.2 -3.5 -0.1 91.8 -5.5 -72.4 -18.7 -21.7 -16.8
+SMC++ (TPAMI’25) [7] -13.1 -4.0 -20.6 -18.1 -39.0 -16.0 -32.2 -3.5 -0.1 91.8 -5.5 -72.4 -18.7 -21.7 -16.8
 SEC-VCM (Ours) -46.2 -44.4 -47.6 -51.2 -64.8 -55.1 -23.6 62.9 -23.0 -44.9 17.9 -86.0 -29.8 -19.2 -38.4
@@
 The proposed SEC-VCM achieves the best rate-task performance, as shown in TABLE II.
-It reduces $46.22\%$ (AP), $44.35\%$ (AP50), and $47.55\%$ (AP75) bitrate compared to VTM.
+It reduces $46.2\%$ (AP), $44.4\%$ (AP50), and $47.6\%$ (AP75) bitrate compared to VTM.
 The detection performance of all compared methods across various bitrates is shown in Fig.
@@
 Figure 10: Visualization on the YTVIS2019 dataset of video instance segmentation results.
-Under intense motion (1st row) and instance overlap (2nd row), Mask2Former [59] achieves more accurate instance contours on the reconstructed video from SEC-VCM.
+Under intense motion (1st row) and instance overlap (2nd row), Mask2Former [55] achieves more accurate instance contours on the reconstructed videos from SEC-VCM.
 IV-E2 Video Instance Segmentation
@@
 As shown in TABLE II, our proposed SEC-VCM outperforms other methods in MOTA and achieves $44.9\%$ bitrate saving compared to VTM.
-Since IDF and IDs are mainly influenced by the inter-frame consistency rather than the frame quality, they show weaker correlation with bitrate.
+Since IDF1 and IDs are mainly influenced by the inter-frame consistency rather than the frame quality, they show weaker correlation with bitrate.
 Although SEC-VCM does not have the best rate-task performance in terms of IDF1 and IDs, it outperforms other methods on variable bitrates, as shown in Fig.
@@
 As demonstrated in TABLE II and Fig.
-11, our proposed SEC-VCM achieves the most bitrate saving compared to VTM in terms of J&F ($29.83\%$) and F ($38.42\%$) while achieving a comparable compression ratio with SMC++ on the metric of J.
+11, our proposed SEC-VCM achieves the most bitrate saving compared to VTM in terms of J&F ($29.8\%$) and F ($38.4\%$) while achieving a comparable compression ratio with SMC++ on the J metric.
 IV-E5 Discussion
@@
 It is worth mentioning that although many methods exhibit excellent rate-distortion (PSNR, MS-SSIM) and rate-perception (LPIPS, etc.) performance that far exceeds standard codecs, their performance in downstream machine vision tasks is not satisfactory.
-These widely used metrics cannot well reflect the needs of machine vision tasks [68], which further illustrates the necessity of VCM-related works.
+These widely used distortion metrics cannot well reflect the needs of machine vision tasks, which further illustrates the necessity of VCM-related works.
+Also, developing general and differentiable metrics tailored for downstream tasks remains an urgent and open problem [64].
 TABLE III: Influence of different backbones on BiEC mechanism. VTM-23.1 is the anchor for calculating BD-rate (%) $\downarrow$. The best rate-task performance is displayed in red.
@@
 VTM-23.1 is the anchor for calculating BD-rate (%) $\downarrow$.
-The best rate-task performance is displayed in red.
+The best rate-task performance is in red.
 Downstream tasks Metric $\lambda_{e}$=$0.1$ $\lambda_{e}$=$0.5$ $\lambda_{e}$=11 $\lambda_{e}$=55 $\lambda_{e}$=$10$
@@
 $M_{6}$ MSE MSE +21.1%
-$M_{7}$ KL KL +17.3%
+$M_{7}$ $\text{KL}_{\text{channel}}$ $\text{KL}_{\text{channel}}$ +17.3%
+$M_{8}$ $\text{KL}_{\text{spatial}}$ $\text{KL}_{\text{spatial}}$ +13.5%
 IV-F2 Ablation on the BiEC Mechanism
@@
 We conduct ablation studies on the bi-directional design and the multi-layer design, as shown in Table V.
-Results from $M_{1}$ and $M_{2}$ demonstrate that each direction of the entropy constraints plays a critical role in improving rate-task performance (average 11.7%).
-Also, the experimental results of $M_{3}$, $M_{4}$, and $M_{5}$ show that removing multi-layer design results in significant rate-task performance degradation (average 12.0%).
-Notably, in $M_{6}$, the proposed conditional entropy loss $\mathcal{L}_{e}$ is replaced with MSE, leading to a 21.1% decrease in rate-task performance, highlighting the superiority of soft alignment over rigid matching.
+Results from $M_{1}$ and $M_{2}$ demonstrate that each direction of the entropy constraints plays a critical role in improving rate-task performance of 11.7%.
+Also, the experimental results of $M_{3}$, $M_{4}$, and $M_{5}$ show that removing multi-layer design results in significant rate-task performance degradation of 12.0%.
+Notably, in $M_{6}$, the proposed conditional entropy loss $\mathcal{L}_{e}$ is replaced by MSE to perform direct element-wise alignment, resulting in a 21.1% decrease in rate-task performance.
+We also applied channel-wise ($M_{7}$) and spatial-wise ($M_{8}$) KL divergence losses to achieve distribution-level alignment, which led to performance drops of 17.3% and 13.5%, respectively.
+These results highlight the superiority of the proposed information-level alignment over direct (MSE) and distribution-level (KL divergence) alignment.
 TABLE VI: Ablation study on the semantic-pixel dual-path fusion.
@@
 Ours 0%
-$M_{8}$ +13.4%
-$M_{9}$ +8.9%
-$M_{10}$ +19.5%
+$M_{9}$ +13.4%
+$M_{10}$ +8.9%
+$M_{11}$ +19.5%
 TABLE VII: Complexity of neural codecs in terms of inference time, encoding time, decoding time, and number of optimized parameters.
@@
 DCVC-RT (CVPR’25) [14] [^6] We turn off its self-implemented CUDA operations and only enable PyTorch-based operations for fair comparison. 81 36 39 20.7M
-PLVC (IJCAI’22) [40] [^7] Since the officially released code of PLVC is not supported by the NVIDIA RTX 3090 and the corresponding CUDA version, it is tested on CPU.
-The discriminator from its GAN architecture is not included in the number of parameters.
+PLVC (IJCAI’22) [37] [^7] Since the officially released code of PLVC is not supported by the NVIDIA RTX 3090 and the corresponding CUDA version, we tested it on CPU.
+Also, the discriminator in its GAN architecture is not included in the number of parameters.
 / / / 19.2M
@@
 SA-ICM (ICIP’24) [8] 720 647 768 76.5M
-SMC++ (ArXiv’24) [6] 1188 / / 96.2M
+SMC++ (TPAMI’25) [7] 1188 / / 96.2M
 SEC-VCM (Ours) 707 584 411 31.9M
-Figure 13: Visualization of the intermediate results in the gated fusion process.
+Figure 13: Visualization results of the gated fusion process.
+Red and blue indicate high and low feature values, respectively.
 The pixel-oriented feature $f_{\textit{pixel}}$ and machine-oriented feature $f_{\textit{semantic}}$ are selectively combined based on $\alpha$ for the final reconstruction.
-Red and blue indicate high and low feature values, respectively.
 IV-F3 Ablation on the SPDF Module
-To explore how to effectively utilize pixel-oriented features $f_{\textit{pixel}}$ for MVS-oriented reconstruction, channel-wise concatenation ($M_{8}$) and direct addition ($M_{9}$) are employed to replace the gated fusion, as shown in Table VI.
+To explore how to effectively utilize pixel-oriented features $f_{\textit{pixel}}$ for MVS-oriented reconstruction, channel-wise concatenation ($M_{9}$) and direct addition ($M_{10}$) are employed to replace the gated fusion, as shown in Table VI.
 Experimental results demonstrate that gated fusion outperforms the other two fusion methods.
-Notably, the rate-task performance drops significantly by $19.5\%$ when $f_{\textit{pixel}}$ is not used ($M_{10}$).
+Notably, the rate-task performance drops significantly by $19.5\%$ when $f_{\textit{pixel}}$ is not used ($M_{11}$).
 Furthermore, visualization results indicate that the gated fusion mechanism can selectively combine $f_{\textit{semantic}}$ and $f_{\textit{pixel}}$ for final reconstruction.
@@
 By aligning the video coding process with the video understanding process of the visual backbone from an entropy perspective, SEC-VCM eliminates machine-vision-irrelevant information.
-The proposed BiEC mechanism softly maintains consistency between the visual backbone’s encoding process and the video decoding process, while the semantic-pixel dual-path fusion module introduces pixel-level priors to suppress artifacts and enhance reconstruction quality.
+The proposed BiEC mechanism softly maintains consistency between the visual backbone’s encoding process and the video decoding process, while the semantic-pixel dual-path fusion module introduces pixel-level priors to suppress artifacts and improve downstream task performance.
 Extensive experiments on diverse downstream tasks demonstrate the superiority and generalization ability of our method in achieving MVS-friendly reconstruction.
-We hope our proposed method offers insights into bridging the gap between neural video coding and machine vision, serving as a step forward for advancing the field of video coding for machines (VCM).
+Nevertheless, our approach still has limitations.
+Developing a more lightweight framework and incorporating multimodal large language models into the downstream task evaluation will be directions for future work.
+We believe our proposed method offers insights into bridging the gap between neural video coding and machine vision, serving as a step forward for advancing the field of video coding for machines.